# LLaVA-MoD: Making LLaVA Tiny via MoE-Knowledge Distillation

### v1

@@ 4.1 Experimental Settings
Implementation Details.
We employ the well-established ”ViT-MLP-LLM” architecture to demonstrate the effectiveness of LLaVA-MoD. Specifically, a pre-trained CLIP-ViT-L/14 is utilized as the vision encoder and a 2-layer MLP is utilized as the adaptor for both the student and the teacher models. Both teacher and student models leverage the Qwen-1.5 and Qwen-2 families to construct their base model. Specifically, the teacher model employs a 7B parameter configuration, while the student model is constructed with 1.8B and 0.5B parameter sizes. The performance of teacher MLLM on common multimodal benchmarks is presented in Table 1.
Table 1: Performance of teacher MLLM on multimodal benchmarks.
@@
As introduced in Section 3.2, the training process comprises mimic distillation and preference distillation, each with specific settings. At initialization, we first freeze the vision encoder and LLM while optimizing the VL adaptor to align image tokens with the word embedding space. This stage employs a cross-entropy loss with a batch size of 512 and a learning rate of 1e-4. In the mimic distillation stage, the vision encoder remains frozen, while the LLM and VL adaptor are co-optimized to distill general knowledge from the teacher MLLM in a dense-to-dense fashion. Then, the FFN in the LLM is first transformed into a sparse architecture with a mixture of FFN experts co-optimized with the VL adaptor to distill specialized knowledge from the teacher MLLM in a dense-to-sparse fashion. This stage employs a KL-divergence loss, with cross-entropy loss added for dense-to-sparse distillation. The batch size is 256, and the learning rate is adjusted to 2e-5. In the preference distillation stage, the model inherits from the mimic distillation. The vision encoder remains frozen, and a mixture of FFN experts is co-optimized with the VL adaptor to distill preference knowledge from the teacher MLLM. This stage employs a DPO loss to optimize the student’s probability for positive responses to be greater than that of the teacher and the student’s probability for negative responses to be lower than that of the teacher. The batch size is 256, and the learning rate is adjusted to 2e-6. Throughout all stages, we employ the adam optimizer (Diederik, 2014) and train on 16 NVIDIA A100 GPUs for one epoch each, totaling approximately 960 GPU hours. The detailed hyperparameters of each training stage are illustrated in Appendix  11.
Training Datasets.
The training data consists of 5M samples from the open-source datasets, with each training stage utilizing a distinct dataset. During Initialization, 0.6M general captioning samples from LLaVA-1.5-pretrain dataset (Liu et al., 2023b) are used to bridge the gap between visual and language modalities. In mimic distillation, 2.4M general captioning and conversation samples are used to distill general knowledge from the teacher MLLM and 1.4M multi-tasks data, including VQA, documents, science, and OCR, are used to distill specialized knowledge from the teacher MLLM. For preference distillation, 0.8M preference data samples are used to transfer preference knowledge from the teacher. The detailed dataset of each training stage is illustrated in Appendix 10.
Evaluation Benchmarks.
@@
4.2 Main Results
In this section, we conduct experiments of LLaVA-MoD to highlight its advantages in two aspects: performance and efficiency. For performance, we evaluate comprehension-oriented benchmarks (Table 2) and hallucination-oriented benchmarks (Table 3). For efficiency, we present a comparison in terms of data samples and model size.
@@ 4.2 Main Results
4.2.1 Comprehension-Oriented Benchmarks
As indicated in Table 2, LLaVA-MoD demonstrates remarkable performance on comprehension-oriented benchmarks. It achieves SOTA average results among the models of 2B size and 1B size. Specifically, The 2B-sized LLaVA-MoD surpasses Mini-Gemini-2B (Li et al., 2024c) by 8.1%, while using a lower image resolution (336 vs. 768). The 1B-sized LLaVA-MoD surpasses SPHINX-Tiny (Gao et al., 2024) by 13.2%, using fewer data samples (5M vs. 15M). Furthermore, LLaVA-MoD-2B matches and even surpasses the performance of large-scale MLLMs. Specifically, The 2B-sized LLaVA-MoD surpasses Qwen-VL-Chat-7B (Bai et al., 2023b) by 8.8% and matches the performance of VILA-3B (Lin et al., 2024b) and MiniCPM-V (Yao et al., 2024). These results highlight that our approach effectively and efficiently training small-scale MLLMs by distilling the sparse MoE architecture.
@@ 4.2.1 Comprehension-Oriented Benchmarks
Model LLM #Param Object POPE MMHal
HalBench Bench
@@
4.2.2 Hallucination-Oriented Benchmarks
As indicated in Table 3, LLaVA-MoD shows remarkable performance in mitigating hallucination, even beating its teacher model. It can be attributed into two aspects: Firstly, by assigning a higher probability for the positive response, preference distillation encourages the student model to focus on providing correct and relevant information. Secondly, by assigning a lower probability for the negative response, preference distillation discourages incorrect or unsubstantiated information. Using the teacher model as a reference to adjust response probabilities, this optimization enables the student model to handle hallucination problems more accurately and reliably, thereby outperforming the teacher model.
Moreover, LLaVA-MoD surpasses similar small-sized models (Yao et al., 2024; Li et al., 2024c; He et al., 2024) and even recent RLHF-based models (Sun et al., 2023; Zhou et al., 2024b; Huang et al., 2024; Leng et al., 2024; Zhou et al., 2023; Zhao et al., 2023b). For instance, on the Object HalBench, it outperforms RLHF-V (Yu et al., 2024a) by 8.2% in response-level hallucination rate and by 21.3% in mention-level hallucination rate. This demonstrates that taking the teacher model as a reference for preference distillation is a challenging task that can encourage the student model to effectively learn to mitigate hallucination.
4.2.3 Efficiency Comparison
We compare the efficiency of LLaVA-MoD with recent SOTA MLLMs in terms of training data and trainable parameters. As indicated in Table 2, The 2B-sized LLaVA-MoD requires only 5M samples and activates merely a subset of its total 2.2B parameters during training and inference. It demonstrates that LLaVA-MoD achieves 8.8% higher performance than Qwen-VL-Chat, using only 23% trainable parameters and 0.3% of training data samples, respectively. Moreover, when compared to the small-sized model MiniCPM-V with 2.8B trainable parameters, LLaVA-MoD also achieves superior performance with just 1.6% of the data samples and 2B parameters, underscoring its efficiency. In conclusion, LLaVA-MoD enables more efficient utilization of data, parameters, and computational resources. This not only results in faster training and inference but also provides a practical solution for deploying high-performance models in resource-constrained environments.
@@ 4.3 Ablation Study
In this section, we begin by analyzing the impact of preference distillation. Next, we conduct ablation studies on a model trained solely with mimic distillation to investigate the effectiveness of our proposed architecture and training strategy on comprehension-oriented benchmarks. We utilize CLIP-ViT-L/14 as the default vision encoder. The default large language models (LLMs) are Qwen-1.5-1.8B and Qwen-1.5-7B for the student and teacher models.
@@ 4.3.1 Impact of Preference Distillation
We extensively evaluated preference distillation, examining its impact on both the comprehension ability and hallucination mitigation capabilities of LLaVA-MoD. As detailed in Table 4 and Table 5, preference distillation does not yield consistent improvements in comprehension across all benchmarks, its impact on hallucination mitigation was strikingly consistent. Across all evaluated benchmarks, preference distillation notably reduced the occurrence of hallucinations, highlighting its efficacy in enhancing the reliability and factual accuracy of LLaVA-MoD. This observation aligns with the field of language modeling (Achiam et al., 2023), have consistently indicated that RLHF techniques tend to prioritize the reduction of hallucinations, often at the expense of potential gains in other areas, such as raw comprehension scores.
Table 4: Performance of preference distillation on comprehension-oriented benchmarks.
@@
Table 5: Performance of preference distillation on hallucination-oriented benchmarks.
Distillation Stage Object POPE MMHal
HalBench Bench
@@ 4.3.2 Impact of Training Strategy
Knowledge Distillation Improves Training Efficiency in MoE.
This study explores the impact of knowledge distillation (KD) and examines how it facilitates MoE model training compared to supervised fine-tuning (SFT). The experiments are conducted using the same sparse configuration of E4T2, where 4 experts are initialized and the top-2 experts are activated. The same datasets and training hyperparameters are employed for both KD and SFT.
Training MoE models is inherently challenging, as it requires effectively leveraging huge data to optimize experts’ routing. Our approach demonstrates that KD, compared to traditional SFT, excels in addressing this challenge, achieving superior performance across multiple benchmarks. As demonstrated in Table 6, KD significantly outperforms SFT on all the benchmarks. Specifically, KD achieves an average improvement of 8.1% over SFT, with particularly pronounced gains in complex multi-task scenarios. For instance, on the MMB and MME benchmarks, KD surpasses SFT by 8.2% and 10.0%, respectively. These results highlight KD’s ability to effectively leverage multiple experts and extract comprehensive knowledge from large-scale MLLMs. By distilling essential knowledge, KD not only maximizes data utilization but also achieves better generalization ability of MoE compared to SFT.
Table 6: Comparison between the KD and SFT. The architecture of the student model is configured as E4T2.
@@
Focusing on Response Improves Distillation Generalization.
This study explores the impact of the distillation target. The default approach distills response, we experiment with distilling both response and instruction. Both settings utilize the same datasets and training hyperparameters. Table 7 shows that distilling only the response achieves superior performance, surpassing jointly distilling response and instruction by 3.1% in terms of average performance. This suggests that focusing on response is more effective for knowledge distillation in auto-regressive modeling. A possible explanation lies in that mimicking instructions could lead to overfitting the specific instruction patterns of training data, potentially reducing the model’s adaptability to unseen instructions. Mimicking the teacher model’s response proves to be effective in improving the student’s auto-regressive modeling generalization capabilities.
Table 7: Comparison between different distillation tokens. The response indicates distilling solely on the generated response tokens, while the response+instruction incorporates additional instruction tokens in the distillation process.
@@ 4.3.3 Impact of Model Architecture
Sparse MoE Enables Effective Knowledge Distillation.
This study explores the effect of sparse architecture in distillation by comparing its performance to dense architecture. We adopt the E4T1 configuration as the sparse architecture to ensure the same activated parameters as the dense architecture. The same datasets and training hyperparameters are used for sparse and dense architectures.
[…]
Table 8: Comparison between sparse and dense architecture within the distillation. The sparse architecture is configured as E4T1.
@@
Teacher Capacity Matters in Knowledge Distillation.
This study explores the impact of the teacher in the distillation process. Specifically, we utilize CLIP-ViT-L/14 as the vision encoder. For the LLM, Qwen-1.5-7B is the strong teacher while Qwen-1.5-4B is the weaker teacher. All experiments are conducted under the same datasets and training hyperparameters to ensure a fair comparison.
As shown in Table 9, a suitable teacher enhances performance across various benchmarks. For instance, the Qwen-1.5-1.8B student model achieves an average performance of 62.3% with the weaker Qwen-1.5-4B teacher. This performance jumps to 63.6% when utilizing the stronger Qwen-1.5-7B teacher. Interestingly, the smaller Qwen-1.5-0.5B student model shows marginal improvement with either the Qwen-1.5-4B or Qwen-1.5-7B teacher. This suggests that an excessively large capacity gap between the teacher and student models can hinder effective knowledge transfer. Utilizing a ”middle teacher” with an intermediate capacity could bridge this gap, facilitating smoother knowledge transfer and boosting the student model’s learning efficiency.
Table 9: Comparison between the strong and weak teachers within the distillation. We set the strong teacher as the LLM is Qwen-1.5-7B and the weak teacher as the LLM is Qwen-1.5-4B. The model configuration is E4T2.
@@ 5 Conclusion
In this paper, we introduce LLaVA-MoD, a novel framework for efficient training of small-scale multimodal language models via knowledge distillation from large-scale ones. It addresses two key challenges in MLLM distillation: enhancing s-MLLM architecture with MoE design for efficiency and expressiveness balance and implementing a progressive knowledge transfer strategy. Extensive experiments show that LLaVA-MoD outperforms existing models with low activation parameters and computational costs. Notably, it outperforms Qwen-VL-Chat-7B by 8.8% with only 2 billion activation parameters, 0.3% training data, and 23% trainable parameters, highlighting its effectiveness in distilling knowledge and driving more efficient multimodal language model development.
6 Limitations
LLaVA-MoD requires that both the teacher and student models belong to the same family of large language models (LLMs) to ensure consistency in the vocabulary space. Future research could explore distillation techniques involving heterogeneous model families. Additionally, the need to load both student and teacher models results in significant memory usage. To facilitate more efficient distillation, a feasible solution would be to pre-extract the outputs of the teacher model, allowing only the student model to be loaded during training.
Appendix A Appendix
A.1 Dataset
We report the detailed training dataset settings of LLaVA-MoD in Table 10.
Table 10: Training dataset. #Sample means the training samples.
Stage Task Dataset
Initialization Captioning LLaVA-1.5-Pretrain (Liu et al., 2023b)
Conversation MIMIC-IT (Li et al., 2023a), LVIS (Wang et al., 2023a), LRV (Liu et al., 2023a), SViT (Zhao et al., 2023a)
Conversation LLaVA-1.5-Instruct (Liu et al., 2023b)
Science AI2D (Kembhavi et al., 2016), ScienceQA (Lu et al., 2022a)
OCR OCRVQA (Mishra et al., 2019), SynthDoG-EN (Kim et al., 2022)
Knowledge A-OKVQA (Schwenk et al., 2022), GeoQA+ (Cao & Xiao, 2022)
Preference Distillation Preference RLAIF-V (Yu et al., 2024b)
A.2 Hyperparameters
We report the detailed training hyperparameters of LLaVA-MoD in Table 11.
Table 11: Training hyperparameters
Configuration Stage-1 Stage-2 Stage-3

### v3

@@ 4.1 Experimental Settings
Implementation Details. We employ the “ViT-MLP-LLM” architecture to demonstrate the effectiveness of LLaVA-MoD. A pre-trained CLIP-ViT-L/14 is utilized as the vision encoder and a 2-layer MLP is utilized as the adaptor. Qwen-1.5/2 with different sizes are utilized as the LLM for l-MLLM and s-MLLM. Specifically, l-MLLM is configured with 7B parameters, while s-MLLM is configured with 1.8B and 0.5B parameters. The performance of l-MLLM on multimodal benchmarks is presented in Tab. 1. We use the same series of LLMs for distillation, i.e. employing Qwen-1.5 7B to distill Qwen-1.5 1.8B and Qwen-1.5 0.5B. Each stage uses distinct training configurations. The detailed training strategy and hyperparameter are illustrated in Appendix A.1.
Table 1: Performance of l-MLLM on multimodal benchmarks.
@@
Training Datasets. The training data consists of 5M samples from the open-source datasets, with each training stage utilizing a distinct dataset. During Initialization, 0.6M general captioning samples are used to bridge the gap between visual and language modalities. In mimic distillation, 2.4M general captioning and conversation samples are used to distill general knowledge from l-MLLM, and 1.4M multi-tasks data, including VQA, documents, science, and OCR, are used to distill specialized knowledge from l-MLLM. For preference distillation, 80K preference data samples are used to transfer preference knowledge from l-MLLM. The detailed dataset of each training stage is illustrated in Appendix A.2.
@@
4.2 Main Results
In this section, we conduct experiments of LLaVA-MoD to highlight its advantages in two aspects: performance and efficiency. For performance, we evaluate comprehension-oriented benchmarks (Tab. 2) and hallucination-oriented benchmarks (Tab. 3). For efficiency, we present a comparison in terms of data samples and model size. The performance is obtained under the same series of LLMs, i.e. employing Qwen-1.5 7B to distill Qwen-1.5 1.8B and Qwen-1.5 0.5B.
Comprehension-Oriented Benchmarks. As indicated in Tab. 2, LLaVA-MoD achieves SOTA average results among the models of 1B and 2B size on comprehension-oriented benchmarks. The 2B-sized LLaVA-MoD surpasses Mini-Gemini-2B (Li et al., 2024) by 8.1%, while using a lower image resolution (336 vs. 768). The 1B-sized LLaVA-MoD surpasses SPHINX-Tiny (Gao et al., 2024) by 13.2%, using fewer data samples (5M vs. 15M). Furthermore, LLaVA-MoD-2B matches and even surpasses the performance of large-scale MLLMs. The 2B-sized LLaVA-MoD surpasses Qwen-VL-Chat-7B (Bai et al., 2023b) by 8.8% and matches the performance of VILA-3B (Lin et al., 2024b) and MiniCPM-V (Yao et al., 2024). These results highlight that our approach trains small-scale MLLMs effectively by distilling the sparse MoE architecture from large-scale MLLMs.
@@ 4.2 Main Results
Model LLM #Param Object HalBench POPE MMHal-Bench
@@
Hallucination-Oriented Benchmarks. As indicated in Tab. 3, LLaVA-MoD shows remarkable performance in mitigating hallucination, even beating its teacher model. It can be attributed to preference distillation from two aspects: Firstly, by assigning a higher probability for the positive response, preference distillation encourages LLaVA-MoD to focus on providing correct and relevant information. Secondly, by assigning a lower probability for the negative response, preference distillation discourages incorrect or unsubstantiated information. By using the teacher model as a reference to adjust response probabilities, preference distillation enables LLaVA-MoD to handle hallucination issues more accurately and reliably, thereby surpassing its teacher model. Moreover, LLaVA-MoD even surpasses recent RLHF-based models (Sun et al., 2023; Zhou et al., 2024b; Huang et al., 2024; Leng et al., 2024; Zhou et al., 2023). On the Object HalBench, it outperforms RLHF-V (Yu et al., 2024a) by 8.2% in response-level hallucination rate and by 21.3% in mention-level hallucination rate. This demonstrates that preference distillation is an effective task to mitigate hallucination.
Efficiency Comparison. We compare the efficiency of LLaVA-MoD with recent SOTA MLLMs in terms of training data and trainable parameters. As indicated in Tab. 2, The 2B-sized LLaVA-MoD requires only 5M samples and activates merely a subset of its total 2.2B parameters during training and inference. It achieves 8.8% higher performance than Qwen-VL-Chat, using only 23% trainable parameters and 0.3% of training data samples. Moreover, compared to the small-sized model MiniCPM-V with 2.8B trainable parameters, LLaVA-MoD also achieves superior performance with just 1.6% of the data samples and 2B parameters, underscoring its efficiency. LLaVA-MoD enables more efficient utilization of data, parameters, and computational resources. This not only results in efficient training and inference but also provides a practical solution for deploying high-performance models in resource-constrained environments.
@@ 4.3 Ablation Study
In this section, we thoroughly explore the effect of the distillation strategy and model design. Firstly, we investigate the preference distillation on comprehension and hallucination benchmarks. Since our focus is on the training strategy and model design, we carry out subsequent ablation experiments specifically on mimic distillation. We use CLIP-ViT-L/14 as the default vision encoder, and Qwen-1.5-1.8B and Qwen-1.5-7B as the default student and teacher LLMs respectively.
@@ 4.3.1 Impact of Preference Distillation
Table 4: The impact of preference distillation on the comprehension-oriented benchmarks.
@@
Table 5: The impact of preference distillation on the hallucination-oriented benchmarks.
Distillation Object HalBench POPE MMHal-Bench
@@
Preference Distillation Mitigates Hallucination. We explore the effect of preference distillation on comprehension capability and hallucination mitigation. As shown in Tab. 4 and Tab. 5, preference distillation remarkably reduces the hallucination of s-MLLM, while it does not yield consistent improvements in comprehension capability. This observation aligns with that in LLMs (Achiam et al., 2023), where preference optimization tends to prioritize reducing hallucinations, often resulting in a certain degree of decline in comprehension gains.
KTO Stabilizes Preference Distillation. We explore the training stability of various preference-optimization losses. We employ KTO (Ethayarajh et al., 2024) and DPO (Rafailov et al., 2024). As shown in Appendix B.2, both methods enhance hallucination mitigation. However, KTO has a significant advantage over DPO. The result implies that KTO, which indicates whether a sample is good or bad, provides a better signal for s-MLLM to surpass l-MLLM compared to DPO, which determines which one is better by comparing two samples.
@@ 4.3.2 Impact of Training Strategy
KD Facilitates MoE Training. We explore the effect of knowledge distillation (KD) in MoE training by comparing it with supervised fine-tuning (SFT). The ablation experiments are carried out using the same sparse configuration of E4T2, where four experts are initialized and the top two experts are activated. As shown in Tab. 6, KD surpasses SFT on all benchmarks, achieving an 8.1% average gain over SFT. Additionally, it achieves notable gains in complex multi-task scenarios, such as a +8.2% gain on MMB and a +10.0% gain on MME. These results indicate that KD provides a better optimization signal for effectively training the MoE architecture.
Table 6: Comparison between KD and SFT. The MoE architecture is E4T2.
@@
General-to-Specialized KD Boosts Performance. We explore the effect of the general-to-specialized approach in mimic distillation by comparing the proposed D2D+D2S with D2S. The D2D+D2S involves a two-phase process. Firstly, it distills a dense s-MLLM using general data. Subsequently, it transforms the s-MLLM from dense to sparse and further distills it using multi-task data. The D2S directly distills the sparse s-MLLM using a combined of general and multi-task dataset. Tab. 7 shows that D2D+D2S surpasses D2S with an average gain of 4.0%. The result indicates that D2D+D2S enables effective knowledge transfer, striking a balance between general and specialized competencies. Additionally, D2D+D2S is more computationally efficient, consuming only 62.5% of the GPU hours required by D2S.
Table 7: Comparison between different distillation strategies in mimic distillation. #GPU indicates the training GPU hours.
[…]
D2S 1536 57.9 36.5 66.3 57.5 65.7 61.4 60.0 57.9
D2D+D2S 960 59.3 40.0 68.4 58.7 68.4 65.1 61.5 60.2
Focus on Response Distillation Improves Generalization. We explores the effect of the distillation target by comparing response distillation with response+instruction distillation. Tab. 8 shows that response distillation surpasses response+instruction distillation with an average gain of 3.1%. The result indicates that focusing on response is more effective for knowledge distillation in auto-regressive modeling. A possible reason for this is that mimicking additional instructions can lead to overfitting the specific instruction patterns in the training data, potentially reducing the generalization to unseen instructions.
Table 8: Comparison between different distilled tokens. Response indicates distilling solely on the response tokens. Response+Instruction incorporates additional distillation of instruction tokens.
@@ 4.3.3 Impact of Model Architecture
Sparse Architecture Facilitates Knowledge Transfer.
We explore the effect of sparse architecture in distillation by comparing it with dense architecture. The configuration is E4T1 where four experts are initialized and the top-1 expert is activated to ensure that the activated parameters are the same as those of the dense architecture. Tab. 9 shows that sparse architecture surpasses dense architecture with an average gain of 3.7%, The superior performance is prominent in complex multi-task benchmarks, such as MME and MMB. The result indicates that sparse architecture leverages MoE to effectively transfer diverse knowledge from l-MLLM while maintaining computational efficiency.
Table 9: Comparison between sparse and dense architecture within the distillation.
@@
Teacher Capacity Matters in Knowledge Distillation. We explore the effect of the teacher capacity in the distillation. We employ Qwen-1.5-7B as the strong teacher and Qwen-1.5-4B as the weaker one. The MoE configuration is set as E4T2. Tab. 10 shows that a well-suited teacher can boost performance. For Qwen-1.5-1.8B, using a 7B-sized teacher yields an average gain of 2.2% compared to a 4B-sized teacher. However, when the student and teacher have a large capacity disparity, it can hinder effective knowledge transfer. For Qwen-1.5-0.5B, using a 7B-sized teacher yields a marginal average gain compared to a 4B-sized teacher. Utilizing a “middle teacher” with an intermediate capacity can bridge the gap, facilitating smooth knowledge transfer.
Table 10: Comparison between the strong and weak teachers. We set the strong teacher as the LLM is Qwen-1.5-7B and the weak teacher as the LLM is Qwen-1.5-4B.
@@ 5 Conclusion
In this paper, we introduce LLaVA-MoD, a novel framework for efficient training of small-scale multimodal language models via knowledge distillation from large-scale ones. It addresses two key challenges in MLLM distillation: enhancing s-MLLM architecture with MoE design for efficiency and expressiveness balance and implementing a progressive knowledge transfer strategy. Extensive experiments show that LLaVA-MoD outperforms existing models with low activation parameters and computational costs. Notably, it outperforms Qwen-VL-Chat-7B by 8.8% with only 2 billion activation parameters, 0.3% training data, and 23% trainable parameters, highlighting its effectiveness in distilling knowledge and driving more efficient MLLM development.
Appendix A Implementation Details
A.1 Training Strategy and Hyperparameters
We first freeze the vision encoder and LLM while optimizing the VL adaptor to align image tokens with the word embedding space during initialization. This stage employs a cross-entropy loss with a batch size of 512 and a learning rate of 1e-4. In mimic distillation, the vision encoder remains frozen, while the LLM and VL adaptor are co-optimized to distill general knowledge from l-MLLM in dense-to-dense distillation. Then, the FFN in the LLM is first transformed into a sparse architecture with a mixture of FFNs co-optimized with the VL adaptor to distill specialized knowledge from l-MLLM in dense-to-sparse distillation. This stage employs a KL-divergence loss, with cross-entropy loss added for dense-to-sparse distillation. The batch size is 256, and the learning rate is adjusted to 2e-5. In preference distillation, the model inherits from the mimic distillation. The vision encoder remains frozen, and a mixture of FFN experts is co-optimized with the VL adaptor to distill preference knowledge from the teacher MLLM. This stage employs a KTO (Ethayarajh et al., 2024) loss to optimize the probability of s-MLLM for positive responses to be greater than that of l-MLLM and the probability of s-MLLM for negative responses to be lower than that of l-MLLM. The batch size is 256, and the learning rate is adjusted to 2e-6. Throughout all stages, we employ the adam optimizer (Diederik, 2014) and train on 16 NVIDIA A100 GPUs for one epoch each, totaling approximately 960 GPU hours. The detailed training hyperparameters are shown in Tab. 11.
Table 11: Training hyperparameters of each stage.
Configuration Initialization Mimic Distillation Preference Distillation